BC Corp. reported EBITDA of $1290 million in 2003, prior to interest expenses of $215 million and depreciation charges of $400 million. Capital Expenditures in 2003 amounted to $450 million, and working capital was 7% of revenues (which were $13,500 million). The firm had debt outstanding of $3.068 billion (in book value terms), trading at a market value of $3.2 billion, and yielding a pre-tax interest rate of 8%. There were 62 million shares outstanding, trading at $64 per share, and the most recent beta is 1.10. The tax rate for the firm is 40%. The Treasury bond rate is 7%. The firm expects revenues, earnings, capital expenditures, and depreciation to grow at 9.5% a year from 2004 to 2008, after which time the growth rate is expected to drop to 4%. (Capital spending will offset depreciation in the steady state period.) The company also plans to lower its debt/equity ratio to 50% for the steady state (which will result in the pre-tax interest rate dropping to 7.5%.) a. Estimate the value of the firm. b. Estimate the value of the equity in the firm and the value per share.
Answer #1

Year

EBITDA

Deprec'n

EBIT

EBIT (1-t)

Cap Exp.

_ WC

FCFF

Term Value

2003

$1,290

$400

$890

$534

$450

$82

$402

2004

$1,413

$438

$975

$585

$493

$90

$440

2005

$1,547

$480

$1,067

$640

$540

$98

$482

2006

$1,694

$525

$1,169

$701

$591

$108

$528

2007

$1,855

$575

$1,280

$768

$647

$118

$578

2008

$2,031

$630

$1,401

$841

$708

$129

$633

$14,326

'2003-2008

After 2008

Cost of Equity =

13.05%

12.30%

AT Cost of Debt =

4.80%

4.50%

Cost of Capital =

(WACC)

9.37%

9.69%

Terminal Value

= {EBIT (1-t)(1+g) - (Rev1998 - Rev1997) * WC as % of Rev}/(WACC-g)

= (841 * 1.04) - (13500 * 1.0955 * 1.04 - 13500 * 1.0955)

* 0.07 /(.0969-.04) = $14,326

A. Value of the Firm= 440/1.0937 + 482/1.09372 + 528/1.09373 + 578/1.09374 + (633 + 14941)/1.09375 = $11,172

B. Value of Equity in the Firm = ($11566 - Market Value of Debt) = 11172 - 3200 = $7,972

Value Per Share = $7,972/62 = $128.57